Problem 11.5. Find the Fourier cosine series of the function f(x): f(x) = 1 +X, 0 < x < .
Answer #1

IIIIIII 1191 11 III LI L IPALVELU f ) = m + 0<x< The Fourier Cosine series of f(x) is given by JT nal n = fe) = 2+ Žan coas (an = [ * ()+1)] - f69-} an = na Therefore the fourier Cosine Series, haq nar og FC) = 3 + Ž 269-1} (as fon) = f(a) = 3x - 62-Feel free to ask any doubt in comment section. Thank you. ?

Please like if you satisfied by my answer. ?